OpenStudy (flexastexas):

c. What is the acceleration due to gravity on the surface of the planet in m/s2? Determine the number of g-forces.

OpenStudy (flexastexas):

3. An astronaut lands on an alien planet. He places a pendulum (L = 0.200 m) on the surface and sets it in simple harmonic motion, as shown in this graph.

OpenStudy (flexastexas):

We also know that it completes 9 cycles in 5 seconds

OpenStudy (flexastexas):

What do you think Mohammed?

OpenStudy (shamim):

U hv more than 1 question in 1 post

OpenStudy (shamim):

Anyway i m trying

OpenStudy (shamim):

Do u hv any idea abt g

OpenStudy (flexastexas):

#3 is not a qustion, it is to let us know the distance

OpenStudy (shamim):

g=gravitational acceleration

OpenStudy (flexastexas):

Yeah, I assume this is the formula we use T = 2pi √(length / local gravity)

OpenStudy (flexastexas):

Correct?

OpenStudy (shamim):

Yes

OpenStudy (shamim):

Both T nd L given

OpenStudy (shamim):

g=?

OpenStudy (flexastexas):

Well that is what I am confused about. What do we use for gravity? Its not earth but how would we determine gravity in this situation?

OpenStudy (shamim):

We can take square of both sides

OpenStudy (flexastexas):

Could you write it out?

OpenStudy (shamim):

Because u hv a square root in ur eqiation

OpenStudy (shamim):

T/(2*pi)=√(L/g)

OpenStudy (shamim):

After squaring both sides we will get (T/(2*pi))^2=L/g

OpenStudy (shamim):

g=?

OpenStudy (flexastexas):

(5/9(2pi)^2?

OpenStudy (flexastexas):

What does squaring both sides do for us?

OpenStudy (shamim):

Squaring cancels root

OpenStudy (shamim):

Anyway i m frm cell phone

OpenStudy (shamim):

So unable to write equation nd drawing

OpenStudy (flexastexas):

Well, whenever you can it would be great. I am going to sleep for now and Ill check back in a while.

OpenStudy (shamim):

After 5 min

OpenStudy (flexastexas):

Ok Ill wait

OpenStudy (shamim):

now i m frm laptop

OpenStudy (shamim):

\[T=2\pi \times \sqrt{L/g}\]

OpenStudy (shamim):

\[\frac{ T }{ 2\pi }=\sqrt{\frac{ L }{ g }}\]

OpenStudy (flexastexas):

0.200/2pi = 0.031830

OpenStudy (shamim):

after squaring we will get \[(\frac{ T }{ 2\pi })^{2}=\frac{ L }{ g }\]

OpenStudy (shamim):

ok

OpenStudy (flexastexas):

Could you explain what information each equation gives us or what direction it takes us

OpenStudy (shamim):

now g=L/0.03183

OpenStudy (flexastexas):

but we need a m/s^2 answer

OpenStudy (shamim):

we can explain the first equation \[T=2\pi \sqrt{\frac{ L }{ g }}\]

OpenStudy (shamim):

anyway L is given in ur question

OpenStudy (shamim):

tell me L=?

OpenStudy (flexastexas):

L= .2 m

OpenStudy (shamim):

correct!!!!!

OpenStudy (shamim):

now put the value of L in ur equation

OpenStudy (shamim):

g=.2/0.03183

OpenStudy (shamim):

g=?

OpenStudy (flexastexas):

6.2844

OpenStudy (flexastexas):

6.2833*

OpenStudy (shamim):

correct !!!!!

OpenStudy (shamim):

thank u !!

OpenStudy (shamim):

u can give me a medal if u r satisfied with my explanation

OpenStudy (flexastexas):

But we need an answer in m/s^2

OpenStudy (flexastexas):

I also found an answer for this question online but want to understand it better. But here is what I found.

OpenStudy (flexastexas):

For small swing angles, the period of an ideal pendulum anywhere is T = 2pi √(length / local gravity) . The astronaut has already done the pendulum and transmitted the data to us, so we can use his data and this formula to calculate the local gravity where he is. P = 2pi √(length / local gravity) 5/9 sec = 2π √(0.2m / gravity) √(0.2m / gravity) = 5/9sec / 2π Take the reciprocal of each side: √(gravity) / √(0.2) = 18π / 5 Multiply each side by √(0.2): √(gravity) = 18π √(0.2) / 5 Square both sides: Gravity = (324 π²) (0.2) / 25 = 25.582 m/sec²

OpenStudy (shamim):

unit of acceleration due to gravity is \[\frac{ m }{ s ^{2} }\]

OpenStudy (shamim):

so the result is different

OpenStudy (shamim):

need to c every steps of our calculation once again

OpenStudy (flexastexas):

I understand this method until his equation turned into √(gravity) / √(0.2) = 18π / 5

OpenStudy (flexastexas):

I am not sure how he got 18π

OpenStudy (shamim):

ok

OpenStudy (shamim):

\[\frac{ 5 }{ 9 } \div 2\pi=\frac{ 5 }{ 9 }\times \frac{ 1 }{ 2\pi }=\frac{ 5 }{ 18\pi }\]

OpenStudy (shamim):

response plz for better understanding

OpenStudy (flexastexas):

Ok, 5/18pi= 0.088419

OpenStudy (flexastexas):

Where do we go from there?

OpenStudy (agent0smith):

(T/(2*pi))^2=L/g If you plug in L=0.2, and T=5/9 into this and then do algebra correctly, you'll have the correct value for g.